# Multiple SLEs and Dyson Brownian motion: transition density and Green’s function

### v2

@@ 2.3 Chordal $N$-$\mathrm{SLE}_{\kappa}$
Let $(\gamma_{1},\ldots,\gamma_{N})\sim\mathbb{P}_{\alpha}^{(\boldsymbol{x})}$ be the chordal $N$-$\mathrm{SLE}_{\kappa}$ in polygon $(\mathbb{H};x_{1},\ldots,x_{2N})$ associated to link pattern $\alpha=\{\{a_{1},b_{1}\},\ldots,\{a_{N},b_{N}\}\}\in\mathrm{LP}_{N}$. We view it as a $2N$-tuple of continuous non-self-crossing curves $\boldsymbol{\eta}=(\eta_{1},\ldots,\eta_{2N})$ as described in Section 1.1: for $j\in\{1,\ldots,N\}$, we define $\eta_{a_{j}}$ to be $\gamma_{j}$ and $\eta_{b_{j}}$ to be the time-reversal of $\gamma_{j}$. In this way, $\eta_{j}$ is a continuous curve in $\mathbb{H}$ starting from $x_{j}$ for $j\in\{1,\ldots,2N\}$. For $j\in\{1,\ldots,2N\}$, if we parameterize $\eta_{j}$ with $a$-usual parameterization, then its driving function $(W^{j}_{t},t\geq 0)$ satisfies (see [PW19, Proposition 4.10] for $\kappa\in(0,4]$ and [FLPW24, Theorem 1.21] for $\kappa\in(4,8)$)
@@ 2.4 Dyson Brownian motion
Fix $\kappa\in(0,8)$ and $a=2/\kappa$. Denote by $p_{*}(t;\boldsymbol{x},\cdot)$ the transition density for Dyson Brownian motion (1.2) with parameter $\beta=8/\kappa$. Then, for any $\boldsymbol{x},\boldsymbol{y}\in\mathfrak{X}_{2N}$,
@@
From the power-law bound (2.4) and (2.5) and the relation (2.1), we know that $f_{3a}(\boldsymbol{x})\mathcal{Z}_{\alpha}(\boldsymbol{x})e^{-\frac{1}{2}|\boldsymbol{x}|^{2}}$ is integrable on $\mathfrak{X}_{2N}$. Thus $\mathcal{J}_{\alpha}$ is a finite constant. ∎
@@
  • $g_{t}^{j}$ is the conformal map from the unbounded connected component of $\mathbb{H}\setminus\eta_{j}([0,t])$ onto $\mathbb{H}$ with $\lim_{z\to\infty}|g_{t}^{j}(z)-z|=0$;
  • $g_{t}$ is the conformal map from the unbounded connected component of $\mathbb{H}\setminus\cup_{j}\eta_{j}([0,t])$ onto $\mathbb{H}$ with $g_{t}(z)=z+2aNt/z+o(1/|z|)$ as $z\to\infty$;
  • $g_{t,j}$ is the conformal map from the unbounded connected component of $\mathbb{H}\setminus\cup_{k\neq j}g_{t}^{j}(\eta_{k}([0,t]))$ onto $\mathbb{H}$ with $\lim_{z\to\infty}|g_{t,j}(z)-z|=0$.
@@ 3.2 $2N$-time local martingales
  • $g_{t_{j}}^{j}$ is the conformal map from the unbounded connected component of $\mathbb{H}\setminus\eta_{j}([0,t_{j}])$ onto $\mathbb{H}$ with $g_{t_{j}}^{j}(z)=z+at_{j}/z+o(1/|z|)$ as $z\to\infty$, $1\leq j\leq 2N$.
  • $g_{\boldsymbol{t}}$ is the conformal map from the unbounded connected component of $\mathbb{H}\setminus\bigcup_{j=1}^{2N}\eta_{j}([0,t_{j}])$ onto $\mathbb{H}$ with $g_{\boldsymbol{t}}(z)=z+\aleph(\boldsymbol{t})/z+o(1/|z|)$ as $z\to\infty$.
  • $g_{\boldsymbol{t},j}$ is the conformal map from the unbounded connecetd component of $\mathbb{H}\setminus\bigcup_{k\neq j}g_{t_{j}}^{j}(\eta_{k}([0,t_{k}]))$ onto $\mathbb{H}$ with $\lim_{z\to\infty}|g_{\boldsymbol{t},j}(z)-z|=0$, $1\leq j\leq 2N$.
@@
Finally, we parameterize these $2N$-tuple of curves with $a$-common parameterization by making the time change $\mathrm{d}t_{j}=g_{\boldsymbol{t},j}^{\prime}(W_{t_{j}}^{j})^{-2}\mathrm{d}t$, then the $2N$-time-parameter local martingale (3.6) becomes $M_{t}^{\alpha}$ in (3.2), and the SDE (3.7) becomes SDE (1.1). Note also that, on the event $\{T>t\}$, the collection $\{\eta_{j}([0,t])\}_{1\leq j\leq 2N}$ has the property that the curve $\eta_{j}([0,t])$ does not disconnect any $\eta_{k}([0,t])$ with $k\neq j$ from $\infty$. This completes the proof. ∎
@@ 4 Gaussian free field level lines and Dyson Brownian motion $\beta=2$
Fix $\kappa\in(0,8)$, $\boldsymbol{x}=(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$, $\alpha\in\mathrm{LP}_{N}$ and suppose $(\gamma_{1},\ldots,\gamma_{N})\sim\mathbb{P}_{\alpha}^{(\boldsymbol{x})}$ is a chordal $N$-$\mathrm{SLE}_{\kappa}$ in $(\mathbb{H};x_{1},\ldots,x_{2N})$ associated to link pattern $\alpha$.
@@
Finally, let us briefly explain how we extend [FWY24, Theorem 1.2] from $\kappa\in(0,4]$ to $\kappa\in(0,8)$ as in Remark 1.5. To prove (1.9), we need the extension to $\kappa\in(4,8)$ of the following three basic inputs:
  • Radon-Nikodym derivative of the law $\mathbb{P}_{\alpha}^{(\mathbb{U};\boldsymbol{x})}$ of chordal $N$-$\mathrm{SLE}_{\kappa}$ against the law of radial Loewner chains driven by $2N$-radial Bessel process with parameter $2a$.
  • Transition density for $2N$-radial Bessel process and its convergence to invariant density as $t\to\infty$ ([HL21, Proposition 5.5]).
  • The quasi-invariant density for the argument process under $\mathbb{P}_{\alpha}^{(\mathbb{U};\boldsymbol{x})}$.
The first input is an analogue of Lemmas 3.1 and 3.2 in this article. Note that the extension to $\kappa\in(4,8)$ of SDE (2.11) for driving function due to recent development in [FPW24, FLPW24] can be modified to the radial case, by using Möbius covariance of pure partition functions and technicals of coordinate changes for $\mathrm{SLE}$ (see [SW05]) respectively. The second input of $2N$-radial Bessel process is an analogue of Lemma 2.6, which is true when parameter $22a\geq 1/$ 2 (i.e., $\kappa\in(0,8]$). The third input is an analogue of Proposition 1.3, and can be now extended to $\kappa\in(4,8)$ due to Lemma 2.3. Based on the extension of these main ingredients, the proof of [FWY24, Theorem 1.2] can be extended to $\kappa\in(4,8)$ using the same analysis.

### v3

@@ 2.2 Chordal $N$-$\mathrm{SLE}_{\kappa}$
Let $(\gamma_{1},\ldots,\gamma_{N})\sim\mathbb{P}_{\alpha}^{(\boldsymbol{x})}$ be the chordal $N$-$\mathrm{SLE}_{\kappa}$ in polygon $(\mathbb{H};x_{1},\ldots,x_{2N})$ associated to link pattern $\alpha=\{\{a_{1},b_{1}\},\ldots,\{a_{N},b_{N}\}\}\in\mathrm{LP}_{N}$. We view it as a $2N$-tuple of continuous non-self-crossing curves $\boldsymbol{\eta}=(\eta_{1},\ldots,\eta_{2N})$ as described in Section 1.1: for $j\in\{1,\ldots,N\}$, we define $\eta_{a_{j}}$ to be $\gamma_{j}$ and $\eta_{b_{j}}$ to be the time-reversal of $\gamma_{j}$. In this way, $\eta_{j}$ is a continuous curve in $\mathbb{H}$ starting from $x_{j}$ for $j\in\{1,\ldots,2N\}$. For $j\in\{1,\ldots,2N\}$, if we parameterize $\eta_{j}$ with $a$-usual parameterization, then its driving function $(W^{j}_{t},t\geq 0)$ satisfies (see [PW19, Proposition 4.10])
@@ 2.3 Dyson Brownian motion
Fix $\kappa\in(0,4]$ and $a=2/\kappa$. Denote by $p_{*}(t;\boldsymbol{x},\cdot)$ the transition density for Dyson Brownian motion (1.2) with parameter $\beta=8/\kappa$. Then, for any $\boldsymbol{x},\boldsymbol{y}\in\mathfrak{X}_{2N}$,
@@
From the power-law bound (2.3) and the relation (2.1), we know that $f_{3a}(\boldsymbol{x})\mathcal{Z}_{\alpha}(\boldsymbol{x})e^{-\frac{1}{2}|\boldsymbol{x}|^{2}}$ is integrable on $\mathfrak{X}_{2N}$. Thus $\mathcal{J}_{\alpha}$ is a finite constant. ∎
@@
  • $g_{t}^{j}$ is the conformal map from $\mathbb{H}\setminus\eta_{j}([0,t])$ onto $\mathbb{H}$ with $\lim_{z\to\infty}|g_{t}^{j}(z)-z|=0$;
  • $g_{t}$ is the conformal map from $\mathbb{H}\setminus\cup_{j}\eta_{j}([0,t])$ onto $\mathbb{H}$ with $g_{t}(z)=z+2aNt/z+o(1/|z|)$ as $z\to\infty$;
  • $g_{t,j}$ is the conformal map from $\mathbb{H}\setminus\cup_{k\neq j}g_{t}^{j}(\eta_{k}([0,t]))$ onto $\mathbb{H}$ with $\lim_{z\to\infty}|g_{t,j}(z)-z|=0$.
@@ 3.2 $2N$-time local martingales
  • $g_{t_{j}}^{j}$ is the conformal map from $\mathbb{H}\setminus\eta_{j}([0,t_{j}])$ onto $\mathbb{H}$ with $g_{t_{j}}^{j}(z)=z+at_{j}/z+o(1/|z|)$ as $z\to\infty$, $1\leq j\leq 2N$.
  • $g_{\boldsymbol{t}}$ is the conformal map from $\mathbb{H}\setminus\bigcup_{j=1}^{2N}\eta_{j}([0,t_{j}])$ onto $\mathbb{H}$ with $g_{\boldsymbol{t}}(z)=z+\aleph(\boldsymbol{t})/z+o(1/|z|)$ as $z\to\infty$.
  • $g_{\boldsymbol{t},j}$ is the conformal map from $\mathbb{H}\setminus\bigcup_{k\neq j}g_{t_{j}}^{j}(\eta_{k}([0,t_{k}]))$ onto $\mathbb{H}$ with $\lim_{z\to\infty}|g_{\boldsymbol{t},j}(z)-z|=0$, $1\leq j\leq 2N$.
@@
Finally, we parameterize these $2N$-tuple of curves with $a$-common parameterization by making the time change $\mathrm{d}t_{j}=g_{\boldsymbol{t},j}^{\prime}(W_{t_{j}}^{j})^{-2}\mathrm{d}t$, then the $2N$-time-parameter local martingale (3.6) becomes $M_{t}^{\alpha}$ in (3.2), and the SDE (3.7) becomes SDE (1.1). This completes the proof. ∎
@@ 4 Gaussian free field level lines and Dyson Brownian motion $\beta=2$
Fix $\kappa\in(0,4]$, $\boldsymbol{x}=(x_{1},\ldots,x_{2N})\in\mathfrak{X}_{2N}$, $\alpha\in\mathrm{LP}_{N}$ and suppose $(\gamma_{1},\ldots,\gamma_{N})\sim\mathbb{P}_{\alpha}^{(\boldsymbol{x})}$ is a chordal $N$-$\mathrm{SLE}_{\kappa}$ in $(\mathbb{H};x_{1},\ldots,x_{2N})$ associated to link pattern $\alpha$.
@@
Finally, let us briefly explain why we can not extend our results in Theorem 1.4 from $\kappa\in(0,4]$ to $\kappa\in(4,8)$. To prove (1.8), we need the extension to $\kappa\in(4,8)$ of the following three basic inputs:
  • Radon-Nikodym derivative of the law $\mathbb{P}_{\alpha}^{(\mathbb{H};\boldsymbol{x})}$ of chordal $N$-$\mathrm{SLE}_{\kappa}$ against the law of chordal Loewner chains driven by Dyson Brownian motion with parameter $\beta=8/\kappa$.
  • Transition density for Dyson Brownian motion and its asymptotic property as $t\to\infty$.
  • Properties of chordal $N$-$\mathrm{SLE}_{\kappa}$ (SDE (2.4) for driving function) and its pure partition functions (power-law bound).
The second input of Dyson Brownian motion is in fact true when parameter $\kappa\in(0,8]$ due to [Rös98]. The third input can also be extended to $\kappa\in(4,8)$ due to recent development in [FPW24, FLPW24]. However, the first input is only valid for $\kappa\in(0,4]$. Indeed, recall that we calculate Randon-Nikodym derivatives by constructing local martingales such as $(M_{t}^{\alpha},0\leq t<T)$ in (3.2). More precisely, we need to define a sequence of stopping times $\tau_{n}:=\inf\{t>0:\mathrm{dist}(\eta_{j}[0,t],\eta_{k}[0,t])<e^{-n},\exists j\neq k\}$ and apply Girsanov’s theorem to martingales $M_{t\wedge\tau_{n}}^{\alpha}$. Note that the exponential part in $M_{t}^{\alpha}$ is related to some Brownian loop measure which is finite only before curves hit each other, i.e., when $t<\tau_{\infty}:=\lim_{n\to\infty}\tau_{n}$. Almost surely, we have $\tau_{\infty}=T$ when $\kappa\in(0,4]$ but $\tau_{\infty}<T$ when $\kappa\in(4,8)$. Therefore, $M_{t}^{\alpha}$ is not well-defined on the whole time interval $t<T$, which is not enough for our subsequent proof.